find the slope of the line that passes through the given points.
Ok are book a little messed up. we are learning about SLopes in Algebra andGeometry
A(5,8) B(7,11)
E(-1,-2), F(-6,-4)
if you could please help me. I need a lot of help.

Answer by Alan3354(69443)

find the slope of the line that passes through the given points.
Ok are book a little messed up. we are learning about SLopes in Algebra andGeometry
Slope = diff of y/diff of x.
Slope is also the rise over the run. For example, a ramp at a loading dock has a slope of the vertical change divided by its horizontal length.
------------------------------
Usually m is used for slope.
----------------
A(5,8) B(7,11)
m = (11-8)/(7-5)
m = 3/2
-------------------------
E(-1,-2), F(-6,-4)
m = (-2 - (-4))/(-1 - (-6)) = (-2 + 4)/(-1 + 6)
m = 2/5
-------------
It makes no difference which point is 1st or 2nd as long as you use the same order for x and y.